# Suppressing spurious transitions using spectrally balanced pulse

### v1

@@ Abstract
Achieving precise control over quantum systems presents a significant challenge, especially in many-body setups, where residual couplings and unintended transitions undermine the accuracy of quantum operations. In superconducting qubits, parasitic interactions—both between distant qubits and with spurious two-level systems—can severely limit the performance of quantum gates. In this work, we introduce a pulse-shaping technique that uses spectrally balanced microwave pulses to suppress undesired transitions. Experimental results demonstrate an order-of-magnitude reduction in spurious excitations between weakly detuned qubits, as well as a substantial decrease in single-qubit gate errors caused by a strongly coupled two-level defect over a broad frequency range. Our method provides a simple yet powerful solution to mitigate adverse effects from parasitic couplings, enhancing the fidelity of quantum operations and expanding feasible frequency allocations for large-scale quantum devices.
@@
Various techniques have been developed to address these challenges [12, 13, 14, 15]. The Derivative Removal by Adiabatic Gate (DRAG) method [16, 17, 18], in particular, was a pulse shaping technique designed to suppress leakage transitions to higher energy states during single-qubit operations. By adding the derivative of the original pulse envelope to the quadrature component, DRAG removes unwanted diabatic leakage transitions, which, in the weak drive regime, show up as a spectral hole at a frequency determined by the prefactor of the derivative. The versatility of DRAG has been investigated in a variety of scenarios, including three-level lambda systems [19], cross-resonance gates [20, 21, 22], and frequency-crowded multilevel systems [23]. However, traditional DRAG corrections face challenges when dealing with weakly detuned transitions, as they struggle to effectively remove spectral components close to the target transition [21].
@@
Figure 1: (a) Schematic diagram illustrating the unintended coupling ($g$) that can occur between two qubits or between a qubit and a spurious TLS. (b) Energy level diagram for the combined system of $\rm Q_{0}$ and $\rm Q_{1}$ (or TLS). Due to state dressing, driving $\rm Q_{0}$ with an amplitude $\Omega$ also cross-drives $\rm Q_{1}$ with a reduced amplitude of ±$\pm g\Omega/\Delta_{0}$ in the weak coupling limit ($g\ll|\Delta_{0}|$). Cross-driving can induce unwanted transitions in $\rm Q_{1}$ when using a pulse with finite width. (c) Time-domain pulse profiles and (d) their normalized Fourier spectra for a 25-ns pulse, showing its original sine-to-the-fourth-power form ($\rm\Omega^{(0)}$), the first-order DRAG-corrected pulse ($\rm\Omega^{(1)}$), and the spectrally balanced pulse with dual-DRAG corrections ($\rm\Omega^{(2)}$). Note that, in the time domain, the real part of $\Omega^{(0)}$ overlaps with that of $\Omega^{(1)}$, while the imaginary part of $\Omega^{(0)}$ overlaps with that of $\Omega^{(2)}$. The vertical dashed lines in (d) indicate the corresponding peak positions. In this case, the frequency to block is assumed to be $\rm\Delta/2\pi=40\ MHz$.
@@
Consider a parasitic exchange-type coupling $g$ between $\rm Q_{0}$ and $\rm Q_{1}$ (or a TLS). Due to the hybridization of the $\left|\rm ge\right\rangle$ and $\left|\rm eg\right\rangle$ states, driving one qubit inevitably induces a cross-driving effect on the other qubit or TLS, but with a reduced amplitude of $g\Omega/\Delta_{0}$, where $\Omega$ is the original drive amplitude applied to $\rm Q_{0}$, and $\Delta_{0}=\omega_{\rm ge}-\omega_{\rm eg}$ is the detuning between the unwanted and target transitions, as illustrated in Fig. 1(b). Natural spectral broadening caused by finite pulse widths can lead to unwanted transitions, such as $\left|\rm gg\right\rangle\leftrightarrow\left|\rm ge\right\rangle$, which degrade the performance of quantum operations. This crosstalk phenomenon is inherently quantum mechanical; the crosstalk Hamiltonian is of the ZX type, exhibiting opposite signs depending on the state of $\rm Q_{0}$. This contrasts with classical signal crosstalk, where the Hamiltonian is of the IX type. As a result, quantum crosstalk cannot be simply corrected by applying a cancellation signal [24].
Interestingly, this quantum crosstalk effect can also be harnessed for beneficial purposes. For instance, the cross-resonance effect, which is achieved by intentionally driving at the resonant frequency of $\rm Q_{1}$, is employed in generating a type of two-qubit entangling gate, the cross-resonance gate [25, 26, 27].
Dual-DRAG — A straightforward solution to suppress an unwanted transition is to remove the spectral components at that transition frequency. A famous example is the DRAG method, which was initially proposed to address leakage transitions to higher energy states in weakly anharmonic systems, such as transmon qubits.
@@
In the DRAG framework, a first-order derivative of the original pulse envelope (which can be complex) is added as a quadrature component. The amplitude factor of the derivative pulse is expressed as $1/\Delta$, where $\Delta$ is the DRAG parameter in units of angular frequency. An example is illustrated in Fig. 1(c). From a frequency-domain perspective, this DRAG correction creates a spectral hole at a frequency $\Delta$ relative to the drive frequency $\omega_{0}$, as depicted in Fig. 1(d). When $\Delta$ is set equal to the qubit anharmonicity $\alpha$, it effectively reduces spectral components near the qubit’s $\left|\rm e\right\rangle$-$\left|\rm f\right\rangle$ transition frequency—typically 200-300 MHz below the $\left|\rm g\right\rangle$-$\left|\rm e\right\rangle$ transition for transmon qubits—thereby suppressing leakage transitions. This technique has also been applied in cross-resonance gates to mitigate off-resonance errors [21] or parasitic transitions [22].
@@
Avoiding weakly detuned transitions during single-qubit operations, which are only a few tens of MHz away, has proven to be challenging [21]. For instance, consider a 25-ns pulse, as shown in Fig. 1(c,d). While a single DRAG correction generates a spectral hole at the frequency defined by $\Delta$, it also shifts the spectral weight to the opposite side. As shown by the example in Fig. 1(d), it leads to an average drive frequency that is about 30 MHz lower than the original drive frequency $\omega_{0}$ given the frequency to block at $\Delta/2\pi=40$ MHz. This effect may be partially mitigated by introducing a constant drive frequency detuning $\eta$ during the pulse [18], i.e. modulating the nominal drive frequency from $\omega_{0}$ to $\omega_{0}+\eta$. However, when the DRAG-induced off-resonance effect is this large, the perturbative treatment in the DRAG theory breaks down. Often, calibrating a high-fidelity gate under such strong drive-induced detuning is experimentally infeasible.
@@
using a set of DRAG parameters $\{\Delta_{1},\Delta_{2},\dots,\Delta_{n}\}$ [28, 22]. Here, $\Omega^{(0)}$ represents the original pulse shape without DRAG. The final pulse $\Omega^{(n)}$ generates multiple spectral holes at the specified frequencies. By two successive DRAG applications at ±$\pm\Delta$, we obtain a pulse envelope that involves a second derivative,
@@
As shown in Fig. 1(d), the spectrum $\hat{\Omega}^{(2)}(\omega)$ remains centered at $\omega_{0}$, while spectral components at $\omega_{0}\pm\Delta$ are eliminated. This dual-DRAG application at mirrored frequencies suppresses weakly detuned transitions and significantly reduces drive-induced off-resonance (see the Supplementary Materials [29] for additional details). For the example presented, we choose the original pulse shape to be a sine raised to the fourth power. This ensures that the first three derivatives vanish at the pulse’s beginning and end, thereby avoiding sharp edges or the need for truncation.
@@
Correcting off-resonance error using virtual-Z compensation — Although the dual-DRAG method substantially reduces the deviation of spectral weighting, system nonidealities such as the presence of higher energy levels and pulse distortions can lead to residual off-resonance effects. To address this, we adopt a post-correction approach that compensates for the overall effect of an off-resonant pulse by appending a VZ gate. We validate this protocol on a device similar to that described in Ref. [30] (see Supplementary Materials [29] for more information). For all experiments in this work unless specified otherwise, we constantly apply one DRAG correction at the qubit anharmonicity $\alpha/2\pi\approx-190$ MHz to prohibit leakage transition to the $\left|\rm f\right\rangle$ state.
@@
Throughout our work, we employ the U3 decomposition rule to synthesize arbitrary single-qubit operations using two $\rm\sqrt{X}$ gates interspersed with VZ gates [31]. Consequently, it suffices to calibrate only the $\rm\sqrt{X}$ gate. It can be shown that any rotation $\rm R$, which maps the qubit from its ground state to the equator (XY plane) of the Bloch sphere, is equivalent to a $\frac{\pi}{2}$ rotation around an axis in the XY plane when combined with a virtual Z or phase rotation (see Supplementary Materials [29] and Ref. [31]). Importantly, since single-qubit gates depend only on relative phase, a $\frac{\pi}{2}$ gate of arbitrary phase can serve as $\rm\sqrt{X}$. Thus, calibrating such an $\rm R$ operation enables the synthesis of any single-qubit gate with the aid of VZ gates.
@@
To calibrate the $\rm R$ gate for a given pulse shape, duration, carrier frequency, and DRAG parameters, we scan the drive amplitude until the excited-state population of the qubit (initialized in its ground state) reaches 0.5 (Fig. 2(a)). By concatenating two $\rm R$ gates with interspersed VZ gates, we then determine the VZ phases required to achieve maximal excitation Fig. 2(b). Based on these initial values, we then fine-tune these parameters using standard pulse train techniques (see supplementary Materials [29] for details). Figure 2(c) shows experimental results for the calibrated parameters $\Omega$ (drive amplitude) and $\phi$ (VZ phase) under intentional drive detunings, which align closely with numerical simulations. Using these VZ-compensated $\rm\sqrt{X}$ gates, we perform randomized benchmarking (RB) [32]. The extracted errors per Clifford (EPC), shown in Fig. 2(d), demonstrate consistent performance across a broad detuning range of over ±$\pm 20$ MHz. Compared to detuning compensation during pulses, the VZ compensation approach is better suited to handle more general and complex scenarios. In practice, when off-resonance effects are strong, balancing the correction of off-resonance errors while avoiding spurious transitions becomes challenging. Our method addresses this challenge effectively.
@@
Figure 3: (a) Experimental setup consisting of two transmon qubits and a tunable coupler, which is also a transmon qubit. The coupler is used to modulate the effective coupling strength between the qubits by adjusting its frequency. (b) Pulse sequence designed for detecting spurious transitions in a nearby qubit during single-qubit gate operations. Repeated $\pi$ gates, each consisting of two $\pi/2$ pulses and spaced by a waiting time $\tau$, are applied to $\rm Q_{0}$. (c) Measured excited-state populations of $\rm Q_{0}$ (top panel) and $\rm Q_{1}$ (bottom panel) using the detection sequence, as a function of the waiting time $\tau$. Results are shown with and without dual-DRAG. Both cases include DRAG correction at the leakage transition to the second excited state $\left|2\right\rangle$. The number of $\pi$ gate pairs is $N=50$. Here, $g/2\pi=1.0~{}\rm{MHz}$, $\Delta_{0}/2\pi=45~{}\rm{MHz}$, and $t_{g}=25~{}\rm{ns}$. (d) Randomized benchmarking results for $\rm Q_{0}$, comparing cases with and without dual-DRAG. (e) Simultaneously monitored excited-state populations of $\rm Q_{1}$. The excitation rates per Clifford gate (ExPC) are $8.4\times 10^{-6}$ with dual-DRAG and $1.6\times 10^{-4}$ without dual-DRAG. (f) Error per Clifford for different coupling strengths, with a fixed detuning of $\Delta_{0}/2\pi=45~{}\rm{MHz}$. (g) Excitation per Clifford for varying coupling strengths. The solid lines are the simulated ExPC averaged over 24 single-qubit Cliffords implemented with the U3 decomposition. The excitation rates are proportional to $g^{2}$ (see Supplementary Materials [29] for details).
@@
Suppressing spurious transitions between qubits — Next, we validate the effectiveness of the dual-DRAG method in suppressing unwanted transitions due to quantum crosstalk. Our investigation is performed on two transmon qubits on the same device. Between them is a tunable coupler—another transmon qubit—which adjusts the effective qubit-qubit coupling strength $g$ via the coupler frequency $\omega_{\rm c}$ (Fig. 3(a)). The qubit frequencies are tunable; their detuning is fixed at $\Delta_{0}/2\pi=45~{}\text{MHz}$. At first, we set $g/2\pi=1$ MHz.
@@
For efficient detection of weak transitions ($g\Omega/\Delta_{0}<\Delta_{0}$), we employ a similar pulse sequence adopted for calibrating the classical microwave crosstalk [33] (Fig. 3(b)). An even number ($2N$) of $\rm X$ gates are applied to the control qubit $\rm Q_{0}$, keeping it in the ground state. Due to non-zero coupling, the spectator qubit $\rm Q_{1}$ is coherently excited by these pulses. A uniform waiting time $\tau$ is added after each $\pi$ gate, during which the wavefunction amplitude of $\rm Q_{1}$’s excited state (e.g., $\left|\rm ge\right\rangle$) accumulates a relative phase $e^{-\mathrm{i}\Delta_{0}\tau}$. At specific $\tau$, the transition amplitudes interfere constructively, amplifying $\rm Q_{1}$’s excitation.
@@
We repeat measurements with dual-DRAG correction at ±$\pm\Delta$. The parameter $\Delta$ is initially set to $\Delta_{0}$ and fine-tuned by minimizing excitation peaks. Using three DRAG parameters $\{\alpha,\Delta,-\Delta\}$, all prominent excitation peaks vanish (Fig. 3(c)), demonstrating effective suppression. Randomized benchmarking on $\rm Q_{0}$ (Fig. 3(d,e)) shows a slight error reduction (error per Clifford: $1.60\times 10^{-3}\to 1.42\times 10^{-3}$). Simultaneously, $\rm Q_{1}$’s excitation rate per Clifford (ExPC) drops by an order of magnitude from $(1.6\pm 0.2)\times 10^{-4}$ to $(0.8\pm 0.3)\times 10^{-5}$. Although the effect in EPC does not seem to be significant, the induced correlated excitations, in particular between distant qubits, complicate the error model and may severely affect error correction codes.
@@
Figure 4: (a) Spectroscopy of a tunable transmon qubit coupled to a spurious TLS, showing an avoided crossing of approximately 10 MHz. The horizontal axis represents the voltage of the flux bias pulse applied to the coupler qubit. (b) Error per Clifford of the qubit as a function of detuning between the qubit and the TLS. Here $t_{g}=25$ ns. (c) Excitation per Clifford of the TLS for different gate times at $\Delta_{0}/2\pi=42~{}\rm{MHz}$. The solid lines are twice the simulated $\rm\sqrt{X}$ gate errors. (d) Excitation per Clifford of the TLS for varying detunings. The solid lines are twice the simulated $\rm\sqrt{X}$ gate errors.
@@
To characterize the dependence on detuning of our method, we calibrate the $\rm\sqrt{X}$ gate at varying qubit-TLS detunings $\Delta_{0}$ using the same protocol developed for coupled qubits. Figure 4(b) shows RB results comparing dual-DRAG corrections (applied at ±$\pm\Delta$) to uncorrected pulses across $20\,\text{MHz}\leq\Delta_{0}/2\pi\leq 85\,\text{MHz}$. For $\Delta_{0}/2\pi<20\,\text{MHz}$, spectral overlap between the drive frequency and the TLS transition hinders the convergence of parameters during calibration, and the dual-DRAG correction overcuts the spectral components around the target frequency, making the drive ineffective. At $\Delta_{0}/2\pi>85\,\text{MHz}$, TLS-induced gate errors diminish to a negligible level.
@@
This approach directly addresses key scalability challenges in a frequency-crowded quantum processor: parasitic qubit couplings and TLS-induced errors, both of which limit frequency allocation. By mitigating these effects, our work can enhance frequency planning flexibility and provide a framework for modeling microwave gate errors in complex environments. The hardware-agnostic nature of the dual-DRAG protocol makes it broadly applicable to other quantum architectures, including quantum dots, NV centers, and trapped ions, positioning it as a versatile tool for high-fidelity quantum control.

### v2

@@ Abstract
Achieving precise control over quantum systems presents a significant challenge, especially in many-body setups, where residual couplings and unintended transitions undermine the accuracy of quantum operations. In superconducting qubits, parasitic interactions—both between distant qubits and with spurious two-level systems—can severely limit the performance of quantum gates. In this work, we introduce a pulse-shaping technique that uses spectrally balanced microwave pulses to suppress undesired transitions. Experimental results demonstrate an order-of-magnitude reduction in spurious excitations between weakly detuned qubits, as well as a substantial decrease in single-qubit gate errors caused by a strongly coupled two-level defect over a broad frequency range. Our method provides a simple yet powerful solution for mitigating adverse effects from parasitic couplings, enhancing quantum gate fidelity. The pulse-shaping technique can be readily adapted to various physical systems.
@@
Various techniques have been developed to address these challenges [5, 12, 30, 43]. The Derivative Removal by Adiabatic Gate (DRAG) method [27, 14, 9, 17], in particular, was a pulse shaping technique designed to suppress leakage transitions to higher energy states during single-qubit operations. By adding the derivative of the original pulse envelope to the quadrature component, DRAG removes unwanted diabatic leakage transitions, which, in the weak drive regime, show up as a spectral hole at a frequency determined by the prefactor of the derivative. The versatility of DRAG has been investigated in a variety of scenarios, including three-level lambda systems [39], cross-resonance gates [24, 41, 21], and frequency-crowded multilevel systems [37]. However, traditional DRAG corrections face challenges when dealing with weakly detuned transitions, as they struggle to effectively remove spectral components close to the target transition [41].
@@
Figure 1: (a) Schematic diagram illustrating the unintended coupling ($g$) that can occur between two qubits or between a qubit and a spurious TLS. (b) Energy level diagram for the combined system of $\rm Q_{0}$ and $\rm Q_{1}$ (or TLS). Due to state dressing, driving $\rm Q_{0}$ with an amplitude $\Omega$ also cross-drives $\rm Q_{1}$ with a reduced amplitude of $\pm g\Omega/\Delta_{0}$ in the weak coupling limit ($g\ll|\Delta_{0}|$). Cross-driving can induce unwanted transitions in $\rm Q_{1}$ when using a pulse with finite width. (c) Time-domain pulse profiles and (d) their normalized Fourier spectra for a 25-ns pulse, showing its original sine-to-the-fourth-power form ($\rm\Omega^{(0)}$), the first-order DRAG-corrected pulse ($\rm\Omega^{(1)}$), and the spectrally balanced pulse with dual-DRAG corrections ($\rm\Omega^{(2)}$). Note that, in the time domain, the real part of $\Omega^{(0)}$ overlaps with that of $\Omega^{(1)}$, while the imaginary part of $\Omega^{(0)}$ overlaps with that of $\Omega^{(2)}$. The vertical dashed lines in (d) indicate the corresponding peak positions. In this case, the frequency to block is assumed to be $\rm\Delta/2\pi=40\,MHz$.
@@
Consider a parasitic exchange-type coupling $g$ between $\rm Q_{0}$ and $\rm Q_{1}$ (or a TLS). Due to the hybridization of the $\left|\rm ge\right\rangle$ and $\left|\rm eg\right\rangle$ states, driving one qubit inevitably induces a cross-driving effect on the other qubit or TLS, but with a reduced amplitude of $g\Omega/\Delta_{0}$, where $\Omega$ is the original drive amplitude applied to $\rm Q_{0}$, and $\Delta_{0}=\omega_{\rm ge}-\omega_{\rm eg}$ is the detuning between the unwanted and target transitions, as illustrated in Fig. 1(b). Natural spectral broadening caused by finite pulse widths can lead to unwanted transitions, such as $\left|\rm gg\right\rangle\leftrightarrow\left|\rm ge\right\rangle$, which degrade the performance of quantum operations. This crosstalk phenomenon is inherently quantum mechanical; the crosstalk Hamiltonian is of the ZX type, exhibiting opposite signs depending on the state of $\rm Q_{0}$. This contrasts with classical signal crosstalk, where the Hamiltonian is of the IX type. As a result, quantum crosstalk cannot be simply corrected by applying a cancellation signal [32]. Interestingly, this crosstalk effect can be leveraged to construct a type of two-qubit entangling gate, the cross-resonance gate [33, 11, 38].
@@
Dual-DRAG — A straightforward solution to suppress an unwanted transition is to remove the spectral components at that transition frequency. A famous example is the DRAG method, which adds a first-order derivative of the pulse envelope as a quadrature component, with an amplitude factor $1/\Delta$, where $\Delta$ is the DRAG parameter in angular frequency units. An example is illustrated in Fig. 1(c). This creates a spectral hole at frequency $\Delta$ relative to the drive frequency, as depicted in Fig. 1(d), effectively reducing spectral components near the qubit’s $\left|\rm e\right\rangle$-$\left|\rm f\right\rangle$ transition — typically 200-300 MHz below the $\left|\rm g\right\rangle$-$\left|\rm e\right\rangle$ transition for transmon qubits — and suppressing leakage transitions.
@@
Avoiding weakly detuned transitions during single-qubit operations, which are only a few tens of MHz away, has proven to be challenging [41]. For instance, as shown in Fig. 1(c,d) a 25-ns pulse with a single DRAG correction of $\Delta/2\pi=40\,{\rm MHz}$ shifts the drive frequency by approximately 30 MHz. When the DRAG-induced off-resonance effect is this large, the perturbative treatment in the DRAG theory breaks down, often causing the conventional mitigation approach which introduces a constant drive frequency detuning $\eta$ [9] to fail.
@@
using a set of DRAG parameters $\{\Delta_{1},\Delta_{2},\dots,\Delta_{n}\}$ [28, 21]. Here, $\Omega^{(0)}$ represents the original pulse shape without DRAG. The final pulse $\Omega^{(n)}$ generates multiple spectral holes at the specified frequencies. By two successive DRAG applications at $\pm\Delta$, we obtain a pulse envelope that involves a second derivative,
@@
As shown in Fig. 1(d), the spectrum $\hat{\Omega}^{(2)}(\omega)$ remains centered at $\omega_{0}$, while spectral components at $\omega_{0}\pm\Delta$ are eliminated. This dual-DRAG application at mirrored frequencies suppresses weakly detuned transitions and significantly reduces drive-induced off-resonance [1]. In the example presented, we choose the original pulse shape to be a sine raised to the fourth power. This ensures that the first three derivatives vanish at the pulse’s beginning and end, thereby avoiding sharp edges or the need for truncation.
Notably, Ref. [17] reports remarkable leakage suppression through their implementation of multiple derivative techniques. Their method, while not explicitly discussed, effectively utilizes an in-phase component analogous to Eq. (2), applying successive DRAG corrections at $\pm\alpha$ followed by an additional correction at $\alpha$. Crucially, this arrangement naturally improves spectral symmetry—a fundamental property that, as we have identified, plays a pivotal role in mitigating DRAG-induced off-resonance effects and enabling their successful high-order derivative suppression.
@@
Correcting off-resonance error using virtual-Z compensation — Although the dual-DRAG method substantially reduces the deviation of spectral weighting, system nonidealities such as the presence of higher energy levels and pulse distortions can lead to residual off-resonance effects. To address this, we adopt a post-correction approach that compensates for the overall effect of an off-resonant pulse by appending a VZ gate. We validate this protocol on a device similar to that described in Ref. [8] (see Supplemental Material [1] for more information). For all experiments in this work unless specified otherwise, we constantly apply one DRAG correction at the qubit anharmonicity $\alpha/2\pi\approx-190\,{\rm MHz}$ to prohibit leakage transition to the $\left|\rm f\right\rangle$ state.
@@
Throughout our work, we employ the U3 decomposition rule to synthesize arbitrary single-qubit operations using two $\rm\sqrt{X}$ gates interspersed with VZ gates [26]. Consequently, it suffices to calibrate only the $\rm\sqrt{X}$ gate. It can be shown that any rotation $\rm R$, which maps the qubit from its ground state to the equator (XY plane) of the Bloch sphere, plus a virtual Z or phase rotation, can be equivalently used as a $\rm\sqrt{X}$ gate (see Supplemental Material [1] and Ref. [26]). Thus, calibrating such an $\rm R$ operation suffices to synthesize any single-qubit gate with the aid of VZ gates.
@@
To calibrate the $\rm R$ gate for a given pulse shape, duration, carrier frequency, and DRAG parameters, we scan the drive amplitude until the excited-state population of the qubit (initialized in its ground state) reaches 0.5 (Fig. 2(a)). We then concatenate two $\rm R$ gates with VZ gates to determine the necessary VZ phases for maximal excitation (Fig.2(b)). After fine-tuning these parameters using standard pulse techniques [1], we validate the calibrated parameters $\Omega$ (drive amplitude) and $\phi$ (VZ phase) under intentional detunings, with results closely matching numerical simulations (Fig. 2(c)). Using these VZ-compensated $\rm\sqrt{X}$ gates, we perform randomized benchmarking (RB) [23]. The errors per Clifford (EPC), shown in Fig. 2(d), demonstrate consistent performance over a detuning range of $\pm 20$ MHz, showing that the VZ compensation method is suited for more general scenarios.
@@
Figure 3: (a) Experimental setup consisting of two transmon qubits and a tunable coupler, which is also a transmon qubit. The coupler is used to modulate the effective coupling strength between the qubits by adjusting its frequency. (b) Pulse sequence designed for detecting spurious transitions in a nearby qubit during single-qubit gate operations. Repeated $\pi$ gates, each consisting of two $\pi/2$ pulses and spaced by a waiting time $\tau$, are applied to $\rm Q_{0}$. (c) Measured excited-state populations of $\rm Q_{0}$ (top panel) and $\rm Q_{1}$ (bottom panel) using the detection sequence, as a function of the waiting time $\tau$. Results are shown with and without dual-DRAG. Both cases include DRAG correction at the leakage transition to the second excited state $\left|2\right\rangle$. The number of $\pi$ gate pairs is $N=50$. Here, $g/2\pi=1\,\rm{MHz}$, $\Delta_{0}/2\pi=45\,\rm{MHz}$, and $t_{g}=25\,\rm{ns}$. (d) Randomized benchmarking results for $\rm Q_{0}$, comparing cases with and without dual-DRAG. (e) Simultaneously monitored excited-state populations of $\rm Q_{1}$. The excitation rates per Clifford gate (ExPC) are $(0.8\pm 0.3)\times 10^{-5}$ with dual-DRAG and $(1.6\pm 0.2)\times 10^{-4}$ without dual-DRAG. (f) Error per Clifford for different coupling strengths, with a fixed detuning of $\Delta_{0}/2\pi=45~\rm{MHz}$. (g) Excitation per Clifford for varying coupling strengths. The solid lines are the simulated ExPC averaged over 24 single-qubit Cliffords implemented with the U3 decomposition. The excitation rates are proportional to $g^{2}$ (see Supplemental Material [1] for details).
@@
Suppressing spurious transitions between qubits — Next, we validate the dual-DRAG method’s effectiveness in suppressing unwanted transitions due to quantum crosstalk. We use two frequency-tunable transmon qubits, with a tunable coupler (another transmon) adjusting the effective qubit-qubit coupling strength $g$ via the coupler frequency $\omega_{\rm c}$ (Fig. 3(a)). At first, we set the qubit-qubit detuning $\Delta_{0}/2\pi=45\,\text{MHz}$ and the coupling $g/2\pi=1\,{\rm MHz}$.
@@
For efficient detection of weak transitions ($g\Omega/\Delta_{0}<\Delta_{0}$), we employ a protocol similar to the interference error filter [22, 3] (Fig. 3(b)). An even number ($2N$) of $\rm X$ gates are applied to the control qubit $\rm Q_{0}$, keeping it in the ground state. Due to non-zero coupling, the spectator qubit $\rm Q_{1}$ is coherently excited by these pulses. A uniform waiting time $\tau$ is added after each $\pi$ gate, during which the wavefunction amplitude of $\rm Q_{1}$’s excited state (e.g., $\left|\rm ge\right\rangle$) accumulates a relative phase $e^{-\mathrm{i}\Delta_{0}\tau}$. At specific $\tau$, the transition amplitudes interfere constructively, amplifying $\rm Q_{1}$’s excitation.
@@
We repeat measurements with dual-DRAG correction at $\pm\Delta$. The parameter $\Delta$ is initially set to $\Delta_{0}$ and fine-tuned by minimizing excitation peaks. Using three DRAG parameters $\{\alpha,\Delta,-\Delta\}$, all prominent excitation peaks vanish (Fig. 3(c)), demonstrating effective suppression. Randomized benchmarking on $\rm Q_{0}$ (Fig. 3(d,e)) shows a slight error reduction (error per Clifford: $1.60\times 10^{-3}\to 1.42\times 10^{-3}$). Simultaneously, $\rm Q_{1}$’s excitation rate per Clifford (ExPC) drops by an order of magnitude from $(1.6\pm 0.2)\times 10^{-4}$ to $(0.8\pm 0.3)\times 10^{-5}$. Although the effect in EPC does not seem to be significant, the induced correlated excitations, in particular between distant qubits, complicate the error model and may severely affect error correction codes.
@@
Figure 4: (a) Spectroscopy of a tunable transmon qubit coupled to a spurious TLS, showing an avoided crossing of approximately 10 MHz. The horizontal axis represents the voltage of the flux bias pulse applied to the coupler qubit. (b) Error per Clifford of the qubit as a function of detuning between the qubit and the TLS. Here $t_{g}=25$ ns. (c) Excitation per Clifford of the TLS for different gate times at $\Delta_{0}/2\pi=42\,\rm{MHz}$. The solid lines are twice the simulated $\rm\sqrt{X}$ gate errors. (d) Excitation per Clifford of the TLS for varying detunings. The solid lines are twice the simulated $\rm\sqrt{X}$ gate errors.
@@
To characterize the dependence on detuning of our method, we calibrate the $\rm\sqrt{X}$ gate at varying qubit-TLS detunings $\Delta_{0}$ using the same protocol developed for coupled qubits. Figure 4(b) shows RB results comparing dual-DRAG corrections (applied at $\pm\Delta$) to uncorrected pulses across $20\,\text{MHz}\leq\Delta_{0}/2\pi\leq 85\,\text{MHz}$. For $\Delta_{0}/2\pi<20\,\text{MHz}$, spectral overlap between the drive frequency and the TLS transition hinders the convergence of parameters during calibration, and the dual-DRAG correction overcuts the spectral components around the target frequency, making the drive ineffective. At $\Delta_{0}/2\pi>85\,\text{MHz}$, TLS-induced gate errors diminish to a negligible level.
@@
This approach directly addresses key challenges in a frequency-crowded quantum processor: parasitic qubit couplings and TLS-induced errors, both of which hinder frequency allocation. By mitigating these effects, our work can enhance frequency planning flexibility and provide a framework for modeling microwave gate errors in complex environments.
We note that while dual-DRAG pulses demonstrate strong performance for small-scale systems (as experimentally validated), their large-scale efficacy requires careful frequency arrangement to avoid extreme conditions with multiple near-resonant TLS (see Supplemental Material [1]). Future work should systematically investigate these scaling properties. However, this scheme can be effectively combined with frequency-tuning approaches that isolate TLS from qubits [10, 7, 40] to further enhance gate performance, providing a comprehensive solution to the TLS problem.
The hardware-agnostic nature of the dual-DRAG protocol makes it broadly applicable to other quantum architectures, including superconducting fluxonium qubits, semiconductor quantum dots, NV centers, and trapped ions, positioning it as a versatile tool for high-fidelity quantum control[16, 13, 36, 39, 35].